# GmNet: Revisiting Gating Mechanisms From A Frequency View

### v2

@@ Abstract
Gating mechanisms have emerged as an effective strategy integrated into model designs beyond recurrent neural networks for addressing long-range dependency problems. In a broad understanding, it provides adaptive control over the information flow while maintaining computational efficiency. However, there is a lack of theoretical analysis on how the gating mechanism works in neural networks. In this paper, inspired by the convolution theorem, we systematically explore the effect of gating mechanisms on the training dynamics of neural networks from a frequency perspective. We investigate the interact between the element-wise product and activation functions in managing the responses to different frequency components. Leveraging these insights, we propose a Gating Mechanism Network (GmNet), a lightweight model designed to efficiently utilize the information of various frequency components. It minimizes the low-frequency bias present in existing lightweight models. GmNet achieves impressive performance in terms of both effectiveness and efficiency in the image classification task.
@@ 1 Introduction
The gating mechanism has been widely used in recurrent neural networks (RNNs) hochreiter1997long ; chung2014empirical to address long-term dependency issues. With the flourish of Transformers vaswani2017attention , self-attention-based architectures have largely replaced RNNs as the core model in natural language processing (NLP). Despite their success, scaling Transformers to long sequences remains challenging due to the quadratic complexity of global attention. Linear gating mechanisms (i.e. Gated Linear Units dauphin2017language ) have been rediscovered as an effective approach to address this problem.
Gated Linear Units (GLUs) introduce a gating mechanism that adaptively controls information flow, allowing the model to focus on relevant features and ignore less important ones. These architectures are simple to implement, and have no apparent computational drawbacks. Therefore, models built with GLUs have been developed in various tasks across NLP de2024griffin ; dubey2024llama ; gu2023mamba and Computer Vision liu2021pay ; ma2024rewrite ; yang2022focal . While architectures based on linear gating mechanisms have demonstrated the effectiveness, their theoretical analysis have not received enough attention. Except intuitive explanations, a deeper examination of their mathematical properties could provide valuable insights for the development of efficient gating mechanisms. Inspired by the Fourier interpretation of element-wise product, we propose to investigate training dynamics of neural networks (NNs) with gating mechanisms from a frequency perspective. This approach helps us understand how frequency components are selectively amplified or suppressed during learning. By the frequency analysis, we aim to uncover how GLUs influence the spectral properties of neural networks and affect the overall performance.
To begin with, we present an intuitive example that illustrates the responses of neural networks to different frequency components of an image. The example as shown in Fig. 1 demonstrates two key points. First, NNs can utilize specific frequency components for classification, not just the low-frequency ones as commonly assumed. This indicates that high-frequency information can also be critical for accurate classification tasks. Second, Gated Linear Units can enhance NNs’ ability of effectively learning and integrating different frequency components.
@@
Figure 1: An illustration of how GLUs affect neural networks in classifying different frequency parts of an image. Starting with a raw image of a frog, we break it down into different frequency bands. The lowest frequency shows a recognizable outline, the middle frequency retains the general shape of the frog, but the highest frequency is almost unrecognizable. Predictions of different components are given in the lower of different models.
In this paper, we systematically explore how gating mechanisms impact the training dynamics. We also provide a comprehensive understanding of how the element-wise product and activation functions interplay to shape the frequency response of neural networks. The first observation is inspired by the convolution theorem and convolution operation in the frequency domain. It indicates that element-wise multiplication in the spatial domain, equivalent to convolution in the frequency domain, inherently introduces new frequency components. It may enhance the learned representations for various frequency domains. However, as demonstrated in prior studies wang2020high ; yin2019fourier , NNs are sensitive to high-frequency noise, which may adversely influence training stability and generalization. Therefore, it is necessary to include activation functions within the gating units to regulate the newly introduced frequency components. Activation functions can broadly be classified based on their differential properties. Smooth activation functions (GELU hendrycks2016gaussian and Swish), characterized by continuous higher-order derivatives, tend to introduce fewer high-frequency components. On the other hand, activation functions (ReLU and ReLU6) with discontinuous higher-order derivatives are more likely to encourage the learning of high-frequency components.
Benefited by the sufficient study from the frequency view, we find that the existing lightweight methods are not aware of how to efficiently and effectively utilize the high-frequency information. Capturing high-frequency components is especially challenging for efficient networks due to their limited capacity and reduced depth. These components often correspond to fine-grained details or subtle semantic distinctions critical for tasks such as fine-grained classification, edge detection, or video understanding. Without adequate modeling of high-frequency information, lightweight models may overly rely on coarse, low-frequency patterns, limiting their generalization and robustness. In this paper, we incorporate the simplest gated linear unit into the current block design and come up with a competitive lightweight model named as Gating Mechanism Network (GmNet). Our proposed network can more effectively capture information across different frequencies, offering both theoretical and practical advantages. With a simplified architecture of the network design, GmNet is able to achieve superior results compared with existing efficient modals without training with advanced techniques. For example, compared to EfficientFormer-L1 li2022efficientformer , GmNet-S3 outperforms $4.0\%$ on top-1 accuracy on ImageNet-1K with 4x faster latency on an A100 GPU.
We succinctly summarize and emphasize the key contributions of this work as following points: 1). We present a comprehensive study on how gated linear units influence neural network training dynamics from a frequency perspective. Through both theoretical analysis and experimental evidences, we demonstrate how the interaction between element-wise multiplication and activation functions shapes the frequency response of NNs. 2).Inspired by the study of the gating mechanism, we investigate the effect of capturing the high-frequency information and propose a competitive lightweight model, named Gating Mechanism Network (GmNet), which achieves promising performance without the need for strong training strategy, surpassing numerous efficient designs.
@@ 2 Related Work
Gated Linear Units. Gating mechanisms have been extensively utilized in neural networks for sequence data processing tasks hochreiter1997long ; gers2000learning ; graves2005framewise ; dauphin2017language . With the rise of Transformer architectures, the Gated Linear Unit shazeer2020glu has been revisited and recognized as an efficient and effective enhancement for modern deep learning models. Recent research has increasingly focused on the integration of gating mechanisms into diverse network architectures of different tasks, including gate-MLP liu2021pay , Mamba gu2023mamba and Llama3 dubey2024llama . While many existing studies provide high-level insights into the functional role of gating mechanisms, there is a notable lack of in-depth analysis of their underlying learning processes. In this work, we conduct a series of experiments combined with solid theoretical analysis to elucidate the roles of gating mechanisms in enhancing the learning of neural networks. Specifically, we investigate the types of information that gating encourages networks to capture from a frequency perspective.
Frequency Learning. Understanding the learning dynamics of neural networks from a frequency perspective is an increasingly interesting topic. It has been found that neural networks respond differently to frequency information depending on the tasks. It has been investigated that neural networks tend to learn low-frequency components quickly before fitting the high-frequency ones for regression tasks yin2019fourier ; tancik2020fourier ; rahaman2019spectral . In classification tasks, by investigating how neural networks respond to different frequency components during training, researchers found that there exists a trade-off between the accuracy and robustness wang2020high . NNs can improve the accuracy by learning more from high frequency components but they are also sensitive to high frequency noise which may disturb the training of NNs. In this paper, we investigate how GLUs optimize the trade-off which drive us to propose a new efficient model design.
Lightweight Networks. The existing lightweight networks can be categorized into two types: (1) Pure convolution-based networks like MobileOne vasu2023mobileone , StarNetma2024rewrite and RepVitwang2024repvit ; (2) Self-attention-based architectures including EfficientModma2024efficient and EfficientFormerV2li2022efficientformer . Although existing network designs already achieve promising performance, they still exhibit significant shortcomings from the perspective of the frequency view. Both convolution and self-attention are learning methods with a low-frequency bias tang2022defects ; bai2022improving . Furthermore, because compact networks typically focus on minimizing parameters and simplifying model structure, their capacity to capture high-frequency details may be restricted. In this paper, we propose a new competitive and efficient network that demonstrates the effectiveness of learning from higher frequency components.
@@ 3 Revisiting Gating Mechanisms from A Frequency View
Figure 2: Block design of different variants of ResNet18 where $\odot$ represents the element-wise product and $\sigma$ means the activation function. In this paper, we investigate GLUs using ReLU6 and GELU as activation functions where corresponding variants can be named as Res18-Gate-ReLU6/GELU.
We begin by defining the components associated with different frequency bands and outlining the details of our experimental setup. With decomposing the raw data $\mathbf{z}$ into the high-frequency part $\mathbf{z}_{h}$ and the low-frequency part $\mathbf{z}_{l}$ where $\mathbf{z}=\mathbf{z}_{h}+\mathbf{z}_{l}$. Denoting a threshold $r$ and an image $\mathbf{x}$, we have the following equations:
$\mathbf{z}=\mathcal{F}(\mathbf{x}),\quad\mathbf{z}_{h},\mathbf{z}_{l}=\theta(\mathbf{z};R)$ (1)
where $\mathbf{z}=\mathcal{F}\left(\mathbf{x}\right)$ is the 2D Discrete Fourier Transform of $\mathbf{x}$. We select three vision backbones including ResNet-18 he2016deep , MobileNetv2 sandler2018mobilenetv2 and EfficientFormer-v2 li2023rethinking as representations to demonstrate the drawbacks of the CNN networks and transformer-based architectures on capturing the high frequency information and how the gating mechanism improve the capability of learning high-frequency components. Modifications to the network blocks of ResNet-18 are depicted in Fig. 2. We evaluate the classification performance on different frequency components of the input images at each training epoch. Changes in accuracy over time provide insights into the learning dynamics within the frequency domain wang2020high . To avoid the occasionality, we calculated the average over three training runs.
Figure 3: Comparison among Res18, Res18-Ewp. Res18-Gate-ReLU6 and Res18-Gate-GELU. We plot the learning curves of Resnet18 and its variants for $100$ epochs, together plotted with the accuracy of different frequency components $\mathbf{z}_{i}$. We set $r$ to $\{4,7,10\}$ respectively. All curves of $\mathbf{z}$ are from the test set. The legends can be found in the top of the figure. More results can be found in the supplementary material.
@@ 3.1 Effect of Element-wise Product
Inspired by the convolution theorem, we first give a theoretical insight of why element-wise product can encourage NNs to learn on various frequency components from a frequency view.
Theorem 1 (Convolution Theorem).
Consider two functions $u(x)$ and $v(x)$ with their Fourier transforms $U$ and $V$:
${(u\cdot v)}(x)=\mathcal{F}^{-1}(U\ast V),$
[…]
@@ 3.2 How Activation Function Works?
We begin with a brief mathematical analysis illustrating how the differential properties of activation functions influence the frequency characteristics of GLUs.
Theorem 2 (Differential Properties of the Fourier Transform).
If a function $f(t)$ is continues in (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), then
$\mathcal{F}[f^{\prime}(t)]=j\omega{F}[\omega]$
Proof.
[…]
The definitions of the symbols are same to Theorem 1.
Lemma 3.
If $f^{(n)}(t)$ is continues in (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), then
$\mathcal{F}[f^{(n)}(t)]=(j\omega)^{n}{F}[\omega],$ (3)
where $f^{(n)}(t)$ represents the $n$-th derivative of $f(t)$. The proof of Lemma 3 can be easily obtained by the mathematical induction based on Theorem 2. From Eq. 3, we can find that $|F(\omega)|\leq\frac{C}{|\omega|^{n}}$ where $C$ is a sufficiently large constant. It indicates that if higher-order derivatives exist and are continuous, the high-frequency components of the Fourier transform decay more quickly with a rate of $|\omega|^{n}$. Conversely, if higher-order derivatives do not exist or are discontinuous, the decay of the high-frequency components of the Fourier transform slows down. Notably, for the smooth activation like $\operatorname{GELU}(x)=x\cdot\frac{1}{2}[1+\operatorname{erf}(x/\sqrt{2})]$, it is infinitely differentiable across the entire real domain, with all orders of derivatives being continuous and finite. The continuity and finite values of higher-order derivatives cause the high-frequency components of the Fourier transform to decay quickly, resulting in fewer high-frequency components.
To validate the above theoretical insights, we conduct experiments on ResNet18 and its variants to examine the impact of gating mechanisms and activation smoothness on frequency learning. As shown in Fig. 3, the testing curves of the model employing non-smooth activation functions (ReLU6) shows visibly surpass above the other models with different thresholds. It demonstrate superior performance in learning from high-frequency components. This supports the hypothesis that reduced smoothness in the activation function enhances the model’s ability to retain high-frequency information. Conversely, the variant using GELU exhibits stronger performance on low-frequency components, reflecting the strength of smooth activations in modeling low-frequency signals.
Figure 4: Comparison among different variants of MobileNetv2 and EfficientFormer-v2 (Eformerv2). Different architectures respond differently to specific frequency component. To ensure an informative comparison, we select representative frequency thresholds tailored to each model where we set $r$ to $\{6,10\}$ for MobileNetv2 and $\{4,8\}$ for EfficientFormer-v2 respectively. Additional results under other threshold configurations are included in the supplementary material.

### v3

@@ Abstract
Lightweight neural networks, essential for on-device applications, often suffer from a low-frequency bias due to their constrained capacity and depth. This limits their ability to capture the fine-grained, high-frequency details (e.g., textures, edges) that are crucial for complex computer vision tasks. To address this fundamental limitation, we perform the first systematic analysis of gating mechanisms from a frequency perspective. Inspired by the convolution theorem, we show how the interplay between element-wise multiplication and non-linear activation functions within Gated Linear Units (GLUs) provides a powerful mechanism to selectively amplify high-frequency signals, thereby enriching the model’s feature representations. Based on these findings, we introduce the Gating Mechanism Network (GmNet), a simple yet highly effective architecture that incorporates our frequency-aware gating principles into a standard lightweight backbone. The efficacy of our approach is remarkable: without relying on complex training strategies or architectural search, GmNet achieves a new state-of-the-art for efficient models. Our code can be found in https://github.com/YFWang1999/GmNet
@@ 1 Introduction
Designing neural networks that are both highly accurate and computationally efficient is a central challenge in modern vision task. Lightweight models are essential for on-device applications, but their reduced capacity often limits their ability to capture the fine-grained details necessary for complex recognition tasks. A growing body of research suggests this limitation stems from a spectral bias, where standard neural network architectures preferentially learn simple, low-frequency global patterns while struggling to capture high-frequency information corresponding to textures and edges Rahaman et al. (2019); Tancik et al. (2020). This fundamental performance gap motivates the exploration of architectural innovations that can improve a model’s representational power without sacrificing efficiency. This bias is particularly pronounced in efficient models whose limited capacity hinders their ability to learn complex, high-frequency information. This limitation motivates our analysis of Gated Linear Units (GLUs)—a computationally inexpensive mechanism already proven effective in various high-performance models De et al. (2024); Liu et al. (2021); Gu and Dao (2023). While their success is often attributed to adaptive information control, their impact on a network’s spectral properties remains largely unexplored. We hypothesize that the element-wise multiplication at the core of GLUs, which corresponds to convolution in the frequency domain, provides a direct mechanism to modulate this spectral bias and enrich a model’s high-frequency learning.
To build intuition, we present an example that visually illustrates how GLUs alter a network’s response to different frequency components of an image as shown in Fig. 1. We take a standard convolution-based lightweight building block (the top one) and create a variant by incorporating our proposed gating unit ( the bottom one). We first provide an input image decomposed into different frequency components from low to high. The visualizations show that the baseline model primarily performs accurate on the low-frequency information, struggling capturing crucial textural details which leads to an incorrect classification on the raw image. In sharp contrast, the model with GLU demonstrates a more balanced spectral response, effectively learning from both low and high-frequency components to form a richer representation. This simple experiment provides initial illustration that gating mechanisms can directly counteract the low-frequency bias in many efficient architectures.
The mechanism enabling this enhanced spectral response is rooted in the convolution theorem: element-wise multiplication in the spatial domain is equivalent to convolution in the frequency domain. This operation allows the network to create complex interactions between different frequency
@@
Figure 1: An illustration of how GLUs affect neural networks in classifying different frequency parts of an image. $\sigma$ means activation function. Starting with a raw image of a ‘Tiger cat’, we break it down into different frequency bands. The lowest frequency shows a recognizable outline, the higher frequency retains the general shape of the cat, but the highest frequency is almost unrecognizable. Predictions of different components are given in the left of different models. This example demonstrates two points: 1. Although low-frequency decomposed images closely resemble the originals, accurate recognition of it does not guarantee accurate recognition of the original images, and 2. GLUs improve the NNs’ ability to learn higher frequency components effectively.
bands, enriching the feature hierarchy. However, as prior work has noted Wang et al. (2020); Yin et al. (2019), naively amplifying high frequencies can make a model overly sensitive to noise. The key, therefore, is selective modulation. We contend that Gated Linear Units, by pairing the multiplication with a data-dependent gate and a non-linear activation function, provide exactly this control. They allow the model to learn when to integrate high-frequency details and how much to trust them, effectively amplifying useful signals while remaining robust to high-frequency noise.
To put these principles into practice, we introduce the Gating Mechanism Network (GmNet), a lightweight architecture designed to leverage the spectral advantages of gating. By effectively capturing information across the full frequency spectrum, GmNet demonstrates that a structurally-motivated design can lead to substantial practical gains. Compared to the existing methods Ma et al. (2024a; b) which also involve gating designs, our design leverages a self-reinforcing gating mechanism in which the modulation and gating signals are derived from a shared representation. This alignment ensures that salient variations, particularly those associated with high-frequency components, are consistently emphasized rather than suppressed. In contrast, methods based on independent projections often act as generic filters, leading to weaker sensitivity to subtle variations that are critical for classification. Consequently, our approach is inherently more effective in preserving and enhancing high-frequency information. The results are compelling: without relying on advanced training techniques, our GmNet-S3 model achieves 81.3% top-1 accuracy on ImageNet-1K. This surpasses EfficientFormer-L1 by a significant 4.0% margin while simultaneously being 4x faster on an A100 GPU, showcasing a new state-of-the-art in efficient network design.
We summarize the key contributions of this work as follows: (1) We provide the first systematic analysis of Gated Linear Units (GLUs) from a frequency perspective, establishing a clear link between their core operations and their ability to modulate a network’s spectral response. (2) We demonstrate that this spectral modulation can directly counteract the inherent low-frequency bias in many lightweight architectures, enabling them to learn more balanced and detailed feature representations from both low and high frequencies. (3) Based on these insights, we introduce the Gating Mechanism Network (GmNet), a simple yet powerful lightweight architecture that achieves a new state-of-the-art in performance and efficiency, validating the practical benefits of our frequency-based design principles.
@@ 2 Related Work
Gated Linear Units. The Gated Linear Unit (GLU) Dauphin et al. (2017), and its modern variants like SwiGLU Shazeer (2020), have become integral components in state-of-the-art deep learning models. Originally developed for sequence processing, their ability to selectively control information flow with minimal computational overhead has led to widespread adoption. In Natural Language Processing, they are central to powerful Transformers such as Llama3 Dubey et al. (2024) and state-space models like Mamba Gu and Dao (2023), where they are lauded for improving training dynamics. This empirical success has spurred their integration into computer vision architectures; models like gMLP Liu et al. (2021) have shown that replacing self-attention with simple gating-MLP blocks can yield competitive performance. However, the prevailing understanding of these mechanisms remains largely functional—they are viewed as adaptive ’information gates.’ A critical gap exists in the analysis of their impact on a network’s fundamental learning properties. Specifically, no prior work has systematically analyzed GLUs from a frequency perspective or connected their operational mechanism to the well-documented problem of low-frequency bias in vision models.
Frequency Learning. Analyzing neural networks from a frequency perspective has revealed a fundamental learning dynamic known as spectral bias: networks of various types consistently learn simple, low-frequency patterns much faster than complex, high-frequency details Rahaman et al. (2019); Yin et al. (2019); Tancik et al. (2020). While initially explored in regression tasks, this bias presents a significant bottleneck for image classification, particularly in lightweight models. Due to their constrained capacity, these models struggle to capture the high-frequency information corresponding to textures and edges, limiting their overall performance. Furthermore, the use of high-frequency components involves a delicate trade-off; while they are critical for accuracy, they can also make models more susceptible to high-frequency noise, impacting robustness Wang et al. (2020). Crucially, while prior work has adeptly characterized these phenomena, it has largely focused on analysis and diagnosis. A clear gap remains in proposing and studying specific, efficient architectural mechanisms that can actively manage this accuracy-robustness trade-off and explicitly counteract spectral bias within a model’s design.
Lightweight Networks. The design of lightweight networks has predominantly followed two streams: pure convolution-based architectures like MobileOne Vasu et al. (2023b) and RepVit Wang et al. (2024), and hybrid approaches incorporating self-attention, such as EfficientFormerV2 Li et al. (2022). While these lines of work have successfully pushed the frontiers of computational efficiency,
Figure 2: Block design of different variants of ResNet18 where $\odot$ represents the element-wise product and $\sigma$ means the activation function.
they are built upon operations that are now understood to have a strong intrinsic low-frequency bias Tang et al. (2022); Bai et al. (2022). This foundational bias is often exacerbated in the lightweight regime; the aggressive optimization for fewer parameters and lower FLOPs further restricts a model’s capacity to learn essential high-frequency information. Consequently, the current paradigm for efficient network design contains a significant blind spot: it has optimized for computational metrics while largely overlooking the spectral fidelity of the learned representations. This leaves a clear opening for new design principles that explicitly aim to correct this low-frequency bias from the ground up.
@@ 3 Revisiting Gating Mechanisms from A Frequency View
We begin by defining the components associated with different frequency bands and outlining the details of our experimental setup. With decomposing the raw data $\mathbf{z}$ into the high-frequency part $\mathbf{z}_{h}$ and the low-frequency part $\mathbf{z}_{l}$ where $\mathbf{z}=\mathbf{z}_{h}+\mathbf{z}_{l}$. We partition the spectrum into two complementary components based on a predefined cutoff radius $r$. Specifically, the frequency representation $z$ is decomposed into a low-frequency part $z_{l}$ and a high-frequency part $z_{h}$, satisfying
$z=z_{l}+z_{h},\quad(z_{l},z_{h})=\Theta(z;r),$ (1)
where $\Theta(\cdot;r)$ represents a radial masking operator that separates spectral coefficients according to their distance from the origin in the frequency plane. In this formulation, coefficients located within radius $r$ are assigned to $z_{l}$, while those outside the radius are assigned to $z_{h}$. This decomposition enables us to analyze model behaviors under controlled spectral partitions.
We select three vision backbones including ResNet-18 He et al. (2016), MobileNetv2 Sandler et al. (2018) and EfficientFormer-v2 Li et al. (2023) as representations to demonstrate the drawbacks of the CNN networks and transformer-based architectures on capturing the high frequency information and how the gating mechanism improve the capability of learning high-frequency components. Modifications to the network blocks of ResNet-18 are depicted in Fig. 2. We evaluate the classification performance on different frequency components of the input images at each training epoch. Changes in accuracy over time provide insights into the learning dynamics within the frequency domain Wang et al. (2020). To avoid the occasionality, we calculated the average over three training runs.
Figure 3: Comparison among Res18, Res18-Ewp, Res18-Gate-ReLU6 and Res18-Gate-GELU. The $r$ represents the threshold of determining the boundary between low-frequency and high-frequency. We plot the learning curves of Resnet18 and its variants for $100$ epochs, together plotted with the accuracy of different frequency components $\mathbf{z}_{i}$. We set $r$ to $10$. All curves of $\mathbf{z}$ are from the test set. The legends can be found in the top of the figure. We also provide more results with different $r$ and different settings in the appendix.
@@ 3.1 Effect of Element-wise Product
Inspired by the convolution theorem, we first give an insight into why element-wise product can encourage NNs to learn on various frequency components from a frequency view. The convolution theorem states that for two functions $u(x)$ and $v(x)$ with Fourier transforms $U$ and $V$,
$(u\cdot v)(x)=\mathcal{F}^{-1}(U\ast V),$ (2)
where $\cdot$ and $\ast$ denote element-wise multiplication and convolution respectively, and $\mathcal{F}$ is the Fourier transform operator defined as $\mathcal{F}[f(t)]=F(\omega)=\int_{-\infty}^{+\infty}f(t)e^{-j\omega t}\,dt$. This indicates that element-wise multiplication in the spatial domain corresponds to convolution in the frequency domain.
To see its implication more clearly, consider the simplest situation: the self-convolution of a function. If the support set of $\mathcal{F}(\omega)$ is $\Omega,\Omega]$, then the support set of $\mathcal{F}\ast\mathcal{F}(\omega)$ will expand to $2\Omega,2\Omega]$. In other words, self-convolution broadens the frequency spectrum. With this enriched frequency content, neural networks have more opportunities to capture and learn from both high-frequency and low-frequency components.
@@ 3.2 How Activation Function Works?
We begin by analyzing how an activation function’s smoothness influences the frequency characteristics of the features it produces. There is a well-established principle in Fourier analysis that connects a function’s smoothness to the decay rate of its Fourier transform’s magnitude. For a function $f(t)$ that is sufficiently smooth (i.e., its $n$-th derivative $f^{(n)}(t)$ exists and is continuous), the magnitude of its Fourier transform, $|F(\omega)|$, is bounded and decays at a rate proportional to $1/|\omega|^{n}$ for large $\omega$. This is a direct consequence of the differentiation property of the Fourier transform:
$\mathcal{F}[f^{(n)}(t)]=(j\omega)^{n}F(\omega),$ (3)
This property implies that the smoother a function is (i.e., the more continuous derivatives it has), the more rapidly its high-frequency components decay.
Conversely, functions with discontinuities or sharp ”corners” where derivatives are undefined (such as the kink in ReLU-like activations) are known to possess significant high-frequency energy. These sharp features require a broad spectrum of high-frequency sinusoids to be accurately represented. This leads to a Fourier transform that decays much more slowly. For example, a function with a simple discontinuity will have a spectrum that decays only at a rate of $1/|\omega|$. Therefore, we hypothesize that non-smooth activation functions will encourage the network to retain and utilize more high-frequency information compared to their smooth counterparts like GELU and Swish, which is infinitely differentiable.
To validate this hypothesis, we conduct an experiment to compare the frequency learning of a representative smooth activation (GELU) against a non-smooth one (ReLU6) within a ResNet18 architecture. As shown in Fig. 3, the model using the non-smooth ReLU6 activation consistently outperforms the GELU variant in learning from high-frequency components across different thresholds. This result supports our hypothesis, illustrating a clear practical difference between these two activation types. The superior performance of ReLU6 on high-frequency data suggests that the slow spectral decay associated with non-smooth functions can be beneficial for tasks requiring fine-grained detail. Conversely, the GELU variant shows a stronger relative performance on low-frequency components, indicating its suitability for capturing broad, structural patterns. While a more exhaustive study is needed, this experiment provides clear evidence for the link between activation smoothness and a model’s spectral learning preferences.
Figure 4: Comparison among different variants of MobileNetv2. Different architectures respond differently to specific frequency component. To ensure an informative comparison, we select representative frequency thresholds tailored to each model where we set $r$ to $10$. Additional results under other threshold configurations and other based models are included in the supplementary material.